# Density Ratio Estimation-based Bayesian Optimization with Semi-Supervised Learning

### v3

@@ 5.1 A Scenario with Unlabeled Point Sampling
We run several synthetic functions for our methods and the baseline methods. For unlabeled point sampling, we sample 100 points from $\mathcal{N}(\mathbf{x}_{1},\mathbf{I}_{d}),\ldots,\mathcal{N}(\mathbf{x}_{n_{l}},\mathbf{I}_{d})$, where $\lfloor n_{u}/n_{l}\rfloor$ or $\lfloor n_{u}/n_{l}\rfloor+1$ points are sampled from each truncated distribution, so that $n_{u}$ points are sampled in total. As shown in Figure 2, our methods outperform the baseline methods incorporating labeled data with unlabeled points. Interestingly, out methods beat GP-based Bayesian optimization. It implies that ours can fairly balance exploration and exploitation. Furthermore, we present the results of learning $\beta$ in Figure 4, where $\beta$ is adaptively selected by (6) every iteration.
(a) Naval
(b) Parkinsons
(c) Slice
@@ 5.2 Scenarios with Fixed-Size Pools
Comparisons of our methods and the existing methods are carried out in these hyperparameter optimization benchmarks (Klein & Hutter, 2019), as reported in Figure 5. We can benchmark a variety of machine learning models, which are defined with specific hyperparameters and trained on one of four datasets: naval propulsion, parkinsons telemonitoring, and slice localization. There exist 62,208 models, which are used as a pool in this paper, for each dataset. Our algorithms show superior performance compared to other approaches. Similar to the synthetic functions, we argue that the adoption of a predefined pool leverages its performance. In some cases, GP-based Bayesian optimization is better than our methods.
@@
(a) Multi-digit MNIST
(b) Results
Figure 7: Examples and results on 64D minimum multi-digit MNIST search. Figure 7(a) shows some examples of multi-digit MNIST and Figure 7(b) demonstrates the results with 20 repeated experiments.
@@ 6 Discussion
We discuss interesting topics on our methods and the properties of DRE-BO-SSL. Moreover, we provide more thorough discussion, limitations, and broader impacts in the appendices.
@@
(a) Beale, LP
(b) Branin, LP
(c) Bukin6, LP
(d) Six-hump camel, LP
(e) Beale, LS
(f) Branin, LS
(g) Bukin6, LS
(h) Six-hump camel, LS
@@ 7 Conclusion
Acknowledgments
This research was supported in part by the University of Pittsburgh Center for Research Computing through the resources provided. Specifically, this work used the H2P cluster, which is supported by NSF award number OAC-2117681.
@@ Appendix A Additional Comparisons of BORE and LFBO
In addition to Figure 1, we include additional comparisons of BORE and LFBO by random forests (Breiman, 2001), gradient boosting (Friedman, 2001), and XGBoost (Chen & Guestrin, 2016) for the Branin function. For Figures 1 and 9, we use $\zeta=0.33$ and $\beta=0.5$.
@@ Appendix D Experiment Details
D.1 Details of Baseline Methods
  • Gaussian process, EI and UCB: It is a Bayesian optimization strategy, which is defined with GP regression with the Matérn 5/2 kernel (Rasmussen & Williams, 2006). Expected improvement (Jones et al., 1998) and GP-UCB (Srinivas et al., 2010) are used as acquisition functions.
  • Random forest, BORE and LFBO: These are BORE and LFBO that employ random forests (Breiman, 2001) with 1000 decision trees. Minimum samples to split are set to 2 for these baselines.
  • Gradient boosting, BORE and LFBO: These methods are BORE and LFBO with gradient boosting classifiers (Friedman, 2001) with 100 decision trees. Learning rate for the classifier is set to 0.3.
  • XGBoost, BORE and LFBO: Similar to gradient boosting, BORE and LFBO with XGBoost have 100 decision trees as base learners with learning rate of 0.3.
  • MLP, BORE and LFBO: These methods are built with two-layer fully-connected networks. The detail of the multi-layer perceptron is described as follows.
The architecture of a multi-layer perceptron is
  1. First layer: fully-connected, input dimensionality $d$, output dimensionality $32$, ReLU;
  2. Second layer: fully-connected, input dimensionality $32$, output dimensionality $1$, Logistic,
where $d$ is the dimensionality of the problem we solve.
Note that most configurations for the baseline methods follow the configurations described in the work (Song et al., 2022).
D.2 Details of Synthetic Benchmarks
@@
D.3 Details of Tabular Benchmarks
@@
D.4 Details of NATS-Bench
@@
D.5 Details of Minimum Multi-Digit MNIST Search
Since multi-digit MNIST, which is composed of images of size $(28,84)$, is high-dimensional, some of the methods used in this work, e.g., methods with random forests, gradient boosting, and XGBoost, struggle to process such data. Therefore, we embed an original image to a lower-dimensional vector using an auxiliary convolutional neural network. The convolutional neural network is trained to classify a three-digit image to one of labels from “000” to “999,” with the following architecture:
@@ 5. Fifth layer: fully-connected, input dimensionality $128$, output dimensionality $64$, ReLU;
The Adam optimizer (Kingma & Ba, 2015) with learning rate $1\times 10^{-3}$ is used to train the network for 100 epochs. To train and test the model fairly, we create a training dataset of 440,000 three-digit images, a validation dataset of 40,000 three-digit images, and a test dataset of 80,000 three-digit images using a training dataset of 55,000 single-digit images, a validation dataset of 5,000 single-digit images, and a test dataset of 10,000 single-digit images in the original MNIST dataset (LeCun et al., 1998). For example, supposing that a test dataset has 1,000 single-digit images per class – it is not true for the MNIST dataset, but it is assumed for explanation – and we would like to generate a three-digit image “753,” $10^{9}$ combinations for “753” can be created. We therefore randomly sample a fixed number of three-digit images from a vast number of possible combinations. In addition, an early stopping technique is utilized by comparing the current validation loss to the average of validation losses for the recent five epochs. Eventually, our network achieves $99.6\%$ in the training dataset, $97.0\%$ in the validation dataset, and $96.9\%$ in the test dataset.
@@ Appendix G Discussion on Pool Sampling
To see the impact of an additional hyperparameter, i.e., the size of a subset of the original pool, which is introduced to speed up semi-supervised learning algorithms, we demonstrate numerical analysis on pool sampling where the size of a predefined pool is 4,000 and $\beta=0.5$ is given. Based on Figures 8 and 14, we can accelerate our framework without significant performance loss.

### v4

@@ 5.1 A Scenario with Unlabeled Point Sampling
We run several synthetic functions for our methods and the baseline methods. For unlabeled point sampling, we sample 100 points from $\mathcal{N}(\mathbf{x}_{1},\mathbf{I}_{d}),\mathcal{N}(\mathbf{x}_{2},\mathbf{I}_{d}),\ldots,\mathcal{N}(\mathbf{x}_{n_{l}},\mathbf{I}_{d})$, where $\lfloor n_{u}/n_{l}\rfloor$ or $\lfloor n_{u}/n_{l}\rfloor+1$ points are sampled from each truncated distribution, so that $n_{u}$ points are sampled in total. As shown in Figure 2, our methods outperform the baseline methods incorporating labeled data with unlabeled points. Interestingly, out methods beat Gaussian process-based Bayesian optimization. It implies that ours can fairly balance exploration and exploitation. Furthermore, we present the results of learning $\beta$ in Figure 4, where $\beta$ is adaptively selected by (9) every iteration.
@@ 5.2 Scenarios with Fixed-Size Pools
Comparisons of our methods and the existing methods are carried out in these hyperparameter optimization benchmarks (Klein & Hutter, 2019), as reported in Figure 5. We can benchmark a variety of machine learning models, which are defined with specific hyperparameters and trained on one of four datasets: naval propulsion, protein structure, Parkinson’s telemonitoring, and slice localization. There exist 62,208 models, which are used as a pool in this paper, for each dataset. Our algorithms show superior performance compared to other approaches. Similar to the synthetic functions, we argue that the adoption of a predefined pool leverages its performance. In some cases, Gaussian process-based Bayesian optimization is better than our methods.
@@
Figure 7: Results with 20 repeated experiments on 64D minimum multi-digit MNIST search.
@@ 6 Discussion
We discuss interesting topics on our methods and the properties of DRE-BO-SSL. Moreover, we provide more thorough discussion and limitations in the appendices.
@@ 7 Conclusion
Acknowledgements
This research was supported in part by the University of Pittsburgh Center for Research Computing through the resources provided. Specifically, this work used the H2P cluster, which is supported by National Science Foundation award number OAC-2117681.
Impact Statement
@@ Appendix A Additional Comparisons of BORE and LFBO
In addition to Figure 1, we include additional comparisons of BORE and LFBO by random forests (Breiman, 2001), gradient boosting (Friedman, 2001), and XGBoost (Chen & Guestrin, 2016) for the Branin function. For Figures 1 and 8, we use $\zeta=0.33$ and $\beta=0.5$.
@@ Appendix D Experiment Details
D.1 Details of Synthetic Benchmarks
@@
D.2 Details of Tabular Benchmarks
@@
D.3 Details of NATS-Bench
@@
D.4 Details of Minimum Multi-Digit MNIST Search
Figure 10: Examples on 64D minimum multi-digit MNIST search.
Since multi-digit MNIST, which is composed of images of size $(28,84)$ and shown in Figure 10, is high-dimensional, some of the methods used in this work, e.g., methods with random forests, gradient boosting, and XGBoost, struggle to process such data. Therefore, we embed an original image to a lower-dimensional vector using an auxiliary convolutional neural network. The convolutional neural network is trained to classify a three-digit image to one of labels from “000” to “999,” with the following architecture:
@@ 5. Fifth layer: fully-connected, input dimensionality $128$, output dimensionality $64$, ReLU;
The Adam optimizer (Kingma & Ba, 2015) with learning rate $1\times 10^{-3}$ is used to train the network for 100 epochs. To train and test the model fairly, we create a training dataset of 440,000 three-digit images, a validation dataset of 40,000 three-digit images, and a test dataset of 80,000 three-digit images using a training dataset of 55,000 single-digit images, a validation dataset of 5,000 single-digit images, and a test dataset of 10,000 single-digit images in the original MNIST dataset (LeCun et al., 1998). For example, supposing that a test dataset has 1,000 single-digit images per class—it is not true for the MNIST dataset, but it is assumed for explanation—and we would like to generate a three-digit image “753,” $10^{9}$ combinations for “753” can be created. We therefore randomly sample a fixed number of three-digit images from a vast number of possible combinations. In addition, an early stopping technique is utilized by comparing the current validation loss to the average of validation losses for the recent five epochs. Eventually, our network achieves $99.6\%$ in the training dataset, $97.0\%$ in the validation dataset, and $96.9\%$ in the test dataset.
@@ Appendix G Discussion on Pool Sampling
(a) Beale, LP
(b) Branin, LP
(c) Bukin6, LP
(d) Six-hump camel, LP
(e) Beale, LS
(f) Branin, LS
(g) Bukin6, LS
(h) Six-hump camel, LS
@@
To see the impact of an additional hyperparameter, i.e., the size of a subset of the original pool, which is introduced to speed up semi-supervised learning algorithms, we demonstrate numerical analysis on pool sampling where the size of a predefined pool is 4,000 and $\beta=0.5$ is given. Based on Figures 14 and 15, we can accelerate our framework without significant performance loss.